The following information is given for n-pentane, C5H12, at 1atm:

boiling point = 36.2 °C Hvap(36.2 °C) = 25.8 kJ/mol
specific heat liquid = 2.28 J/g°C


At a pressure of 1 atm, what is H in kJ for the process of condensing a 21.0 g sample of gaseous n-pentane at its normal boiling point of 36.2 °C.
1
Expert's answer
2019-07-31T12:55:46-0400

The following information is given for n-pentane, C5H12, at 1atm:

boiling point = 36.2 °C Hvap(36.2 °C) = 25.8 kJ/mol

specific heat liquid = 2.28 J/g°C

At a pressure of 1 atm, what is H in kJ for the process of condensing a 21.0 g sample of gaseous n-pentane at its normal boiling point of 36.2 °C.

Solution

Hcon=-Hvap=-25,8 kJ/mol

H=Hcon*n(C5H12)

n(C5H12)=m/M(C5H12)=21,0/72,2=0,29 mol

H=-25,8*0,29=-7,5 kJ

Answer H=-7,5 kJ


P.S. fixed